OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x^2-7x-6=0

OpenStudy (anonymous):

\[- b \pm \sqrt{\frac{ b^2-4ac }{ 2a }}\] a=1, b=-7,c=-6. \[-(-7) \pm \sqrt{\frac{ (-7^2-4(1)(-6) }{ 2(1) }}\] \[7 \pm \sqrt{\frac{ 49+24 }{ 2 }}\] \[7 \pm \sqrt{\frac{73}{2}}\] \[x=7 + \sqrt{26.5}\] and \[x=7-\sqrt{26.5}\]

OpenStudy (anonymous):

The way you put it is confusing... Can you please type it out better? @halorazer